A ballon is filled with helium, and it’s volume is 2.2L at 298 K and at atmospheric pressure. the ballon bursts if it’s volume exceeds 2.3L. At what temperature would you expect the ballon to burst if the atmospheric pressure remains a constant
1
Expert's answer
2020-04-19T07:37:56-0400

T1=298 K

V1=2.2 L

If the ballon bursts the final: - volume: V2=2.3 L

-Temperature: T2=x

V1/T1=V2/T2;

T2=T1*V2/V1=298*2.3/2.2=311.5 K

Answer: 311.5 K